# Learn2Mix: Training Neural Networks Using Adaptive Data Integration

### v2

@@ B.3 CIFAR-10 Dataset
The test dataset, $K$, comprises $N_{\text{test}}=10000$ samples, with class proportions equivalent to the class proportions in the base CIFAR-10 test dataset. For CIFAR-10 reconstruction (see Section 4.3), we use manual class imbalancing, reducing the number of samples in numerical classes $1-4,7-10$ by a factor of 10. The original training dataset, $J$, now has $N=14000$ samples. The fixed-proportion mixing parameters (for default numerical class ordering of categories from $1-10$) are:
@@
B.4 Imagenette Dataset
The Imagenette dataset contains a subset of $10$ classes from the ImageNet dataset of color images, and is commonly used to train image processing systems. The base Imagenette training dataset, $I$, comprises $N_{I}=9469$ samples, and the base Imagenette test dataset, $K$, comprises $N_{\text{test}}=3925$ samples. For the Imagenette classification result in Section 4.1, we utilize manual class imbalancing. Let $N_{i}\in\mathbb{N}$ be the number of samples in each class, $i\in\{1,\ldots,10\}$, from $I$, where $\smash{N_{I}=\sum_{i=1}^{10}N_{i}}$. We define $\epsilon_{i}=1-0.1i$, $\forall i\in\{1,\ldots,10\}$ as the linearly decreasing imbalance factor. Accordingly, the original training dataset, $J$, has $\smash{N=\sum_{i=1}^{10}\epsilon_{i}N_{i}=5207}$ samples. The fixed-proportion mixing parameters (for default numerical class ordering of categories from $1-10$) are:
@@ B.4 Imagenette Dataset
B.5 CIFAR-100 Dataset
The CIFAR-100 dataset is a collection of color images categorized into $100$ different classes, and is commonly used to train image processing systems. The base CIFAR-100 training dataset, $I$, has $N_{I}=50000$ samples, and the base CIFAR-100 test dataset, $K$, has $N_{\text{test}}=10000$ samples. For the CIFAR-100 classification result in Section 4.1, we utilize manual class imbalancing. Let $N_{i}\in\mathbb{N}$ be the number of samples in each class, $i\in\{1,\ldots,100\}$, from $I$, whereby $\smash{N_{I}=\sum_{i=1}^{100}N_{i}}$. We define $\smash{\epsilon_{i}=40^{-i/100}}$, $\forall i\in\{1,\ldots,100\}$ as the logarithmically decreasing imbalance factor. Accordingly, the original training dataset, $J$, has $\smash{N=\sum_{i=1}^{100}\epsilon_{i}N_{i}=13209}$ samples. The fixed-proportion mixing parameters (for default numerical class ordering of categories from $1-100$) are:
@@ B.6 IMDB Dataset
B.7 Mean Estimation Dataset
The Mean Estimation dataset is a synthetic benchmark designed for regression tasks, wherein each example, $(x_{j},y_{j})$, comprises a 10-dimensional feature vector, $x_{j}$, of samples from one of four statistical distributions, and the mean, $y_{j}$, of this distribution. We create an imbalanced original training dataset, $J$, with $N=3000$ samples, where $J_{1}$ has $1000$ examples drawn from a normal distribution with $\sigma=1$, $J_{2}$ has $1000$ examples drawn from an exponential distribution, $J_{3}$ has $800$ examples drawn from a chi-squared distribution, and $J_{4}$ has $200$ samples drawn from a uniform distribution. The fixed-proportion mixing parameters (for numerical ordering of distributions from $1-4$) are:
@@ B.7 Mean Estimation Dataset
B.8 Wine Quality Dataset
The Wine Quality dataset consists of physicochemical tests on white and red wine samples, and the corresponding quality rating. We treat the wine type (white $=1$, red $=2$) as a categorical variable, wherein $k=2$. We split the Wine Quality dataset such that the base Wine Quality training dataset, $J$, has $N=3248$ samples, and the base Wine Quality test dataset, $K$, has $N_{\text{test}}=3249$ samples. For the Wine Quality regression result in Section 4.2, we utilize manual class imbalancing, reducing the number of samples in numerical class $1$ by a factor of 10. The original training dataset, $J$, now has $N=1043$ samples, where the fixed-proportion mixing parameters (for numerical class ordering of wine type from $1,2$) are:
@@ B.8 Wine Quality Dataset
B.9 California Housing Dataset
The California Housing dataset contains housing data from California and their associated prices. As the ocean proximity variable is categorical ($<$1H OCEAN $=1$, INLAND $=2$, NEAR BAY $=3$, NEAR OCEAN $=4$), we denote $k=4$. We split the California Housing dataset such that the base California Housing training dataset, $J$, has $N=10214$ samples, and the base California Housing test dataset, $K$, has $N_{\text{test}}=10214$ samples. For the California Housing regression result in Section 4.2, we use manual class imbalancing, reducing the number of samples in numerical classes $1,2,4$ by a factor of 20. The original training dataset, $J$, now has $N=3641$ samples. The fixed-proportion mixing parameters (for numerical class ordering of ocean proximity from $1-4$) are:
@@ B.9 California Housing Dataset
Appendix C Experiment Details
C.1 Neural Network Architectures
@@
C.1.1 Fully Connected Networks
@@
C.1.2 Convolutional Neural Networks
@@
For image classification on CIFAR-10 and CIFAR-100, we employ an adapted, larger version of the LeNet-5 model. The network consists of the following layers, wherein $k=10$ for CIFAR-10 and $k=100$ for CIFAR-100.
@@
C.1.3 Residual Neural Networks
@@ C.3 SMOTE Training
C.4 Neural Network Training Hyperparameters
The relevant hyperparameters used to train the neural networks outlined in Section C.1 are given in Table 3. All results presented in the main text were produced using these hyperparameter choices.
@@ C.4 Neural Network Training Hyperparameters
Dataset Task Optimizer Learning Rate ($\boldsymbol{\eta}$) Mixing Rate ($\boldsymbol{\gamma}$) (Learn2Mix) Batch Size ($\boldsymbol{M}$)
MNIST Classification Adam 1.0e-5 0.1 1000
Fashion MNIST Classification Adam 5.0e-6 0.5 1000
CIFAR-10 Classification Adam 1.0e-5 0.1 1000

### v3

@@ C.3 CIFAR-10 Dataset
The test dataset, $K$, comprises $N_{\text{test}}=10000$ samples, with class proportions equivalent to the class proportions in the base CIFAR-10 test dataset. For CIFAR-10 reconstruction (see Section 4.3), we use manual class imbalancing, reducing the number of samples in numerical classes 14,7101-4,7-10 by a factor of 10. The original training dataset, $J$, now has $N=14000$ samples. The fixed-proportion mixing parameters (for default numerical class ordering of categories from 1101-10) are:
@@
C.4 Imagenette Dataset
The Imagenette dataset contains a subset of $10$ classes from the ImageNet dataset of color images, and is commonly used to train image processing systems. The base Imagenette training dataset, $I$, comprises $N_{I}=9469$ samples, and the base Imagenette test dataset, $K$, comprises $N_{\text{test}}=3925$ samples. For the Imagenette classification result in Section 4.1, we utilize manual class imbalancing. Let $N_{i}\in\mathbb{N}$ be the number of samples in each class, $i\in\{1,\ldots,10\}$, from $I$, where $\smash{N_{I}=\sum_{i=1}^{10}N_{i}}$. We define $\epsilon_{i}=1-0.1i$, $\forall i\in\{1,\ldots,10\}$ as the linearly decreasing imbalance factor. Accordingly, the original training dataset, $J$, has $\smash{N=\sum_{i=1}^{10}\epsilon_{i}N_{i}=5207}$ samples. The fixed-proportion mixing parameters (for default numerical class ordering of categories from 1101-10) are:
@@ C.4 Imagenette Dataset
C.5 CIFAR-100 Dataset
The CIFAR-100 dataset is a collection of color images categorized into $100$ different classes, and is commonly used to train image processing systems. The base CIFAR-100 training dataset, $I$, has $N_{I}=50000$ samples, and the base CIFAR-100 test dataset, $K$, has $N_{\text{test}}=10000$ samples. For the CIFAR-100 classification result in Section 4.1, we utilize manual class imbalancing. Let $N_{i}\in\mathbb{N}$ be the number of samples in each class, $i\in\{1,\ldots,100\}$, from $I$, whereby $\smash{N_{I}=\sum_{i=1}^{100}N_{i}}$. We define $\smash{\epsilon_{i}=40^{-i/100}}$, $\forall i\in\{1,\ldots,100\}$ as the logarithmically decreasing imbalance factor. Accordingly, the original training dataset, $J$, has $\smash{N=\sum_{i=1}^{100}\epsilon_{i}N_{i}=13209}$ samples. The fixed-proportion mixing parameters (for default numerical class ordering of categories from 11001-100) are:
@@ C.6 IMDB Dataset
C.7 Mean Estimation Dataset
The Mean Estimation dataset is a synthetic benchmark designed for regression tasks, wherein each example, $(x_{j},y_{j})$, comprises a 10-dimensional feature vector, $x_{j}$, of samples from one of four statistical distributions, and the mean, $y_{j}$, of this distribution. We create an imbalanced original training dataset, $J$, with $N=3000$ samples, where $J_{1}$ has $1000$ examples drawn from a normal distribution with $\sigma=1$, $J_{2}$ has $1000$ examples drawn from an exponential distribution, $J_{3}$ has $800$ examples drawn from a chi-squared distribution, and $J_{4}$ has $200$ samples drawn from a uniform distribution. The fixed-proportion mixing parameters (for numerical ordering of distributions from 141-4) are:
@@ C.7 Mean Estimation Dataset
C.8 Wine Quality Dataset
The Wine Quality dataset consists of physicochemical tests on white and red wine samples, and the corresponding quality rating. We treat the wine type (white =1=1, red =2=2) as a categorical variable, wherein $k=2$. We split the Wine Quality dataset such that the base Wine Quality training dataset, $J$, has $N=3248$ samples, and the base Wine Quality test dataset, $K$, has $N_{\text{test}}=3249$ samples. For the Wine Quality regression result in Section 4.2, we utilize manual class imbalancing, reducing the number of samples in numerical class $1$ by a factor of 10. The original training dataset, $J$, now has $N=1043$ samples, where the fixed-proportion mixing parameters (for numerical class ordering of wine type from $1,2$) are:
@@ C.8 Wine Quality Dataset
C.9 California Housing Dataset
The California Housing dataset contains housing data from California and their associated prices. As the ocean proximity variable is categorical (<<1H OCEAN =1=1, INLAND =2=2, NEAR BAY =3=3, NEAR OCEAN =4=4), we denote $k=4$. We split the California Housing dataset such that the base California Housing training dataset, $J$, has $N=10214$ samples, and the base California Housing test dataset, $K$, has $N_{\text{test}}=10214$ samples. For the California Housing regression result in Section 4.2, we use manual class imbalancing, reducing the number of samples in numerical classes $1,2,4$ by a factor of 20. The original training dataset, $J$, now has $N=3641$ samples. The fixed-proportion mixing parameters (for numerical class ordering of ocean proximity from 141-4) are:
@@ C.9 California Housing Dataset
Appendix D Experiment Details
D.1 Neural Network Architectures
@@
D.1.1 Fully Connected Networks
@@
D.1.2 Convolutional Neural Networks
@@
For image classification on CIFAR-10 and CIFAR-100, we employ an adapted, larger version of the LeNet-5 model, which we call ‘Large LeNet’. The network consists of the following layers, wherein $k=10$ for CIFAR-10 and $k=100$ for CIFAR-100.
@@
D.2 Mobile Neural Networks
For image classification on CIFAR-10 and CIFAR-100, we also employ the MobileNet-V3 Small architecture. The network consists of the following layers, where $k=10$ for CIFAR-10 and $k=100$ for CIFAR-100.
  • Convolutional Stem (features0): 3 input channels, 16 output channels, kernel size 3, stride 2, padding 1, followed by BatchNorm and Hard-Swish activation.
  • Inverted Residual Block 1 (features1): expansion factor 1, 16 to 16 channels, kernel size 3, stride 2, SE disabled, activation ReLU.
  • Inverted Residual Block 2 (features2): expansion factor 4.5, 16 to 24 channels, kernel size 3, stride 2, SE disabled, activation ReLU.
  • Inverted Residual Block 3 (features3): expansion factor 3.67, 24 to 24 channels, kernel size 3, stride 1, SE disabled, activation ReLU.
  • Inverted Residual Block 4 (features4): expansion factor 4, 24 to 40 channels, kernel size 5, stride 2, SE enabled, activation Hard-Swish.
  • Inverted Residual Block 5 (features5): expansion factor 6, 40 to 40 channels, kernel size 5, stride 1, SE enabled, activation Hard-Swish.
  • Inverted Residual Block 6 (features6): expansion factor 6, 40 to 40 channels, kernel size 5, stride 1, SE enabled, activation Hard-Swish.
  • Inverted Residual Block 7 (features7): expansion factor 3, 40 to 48 channels, kernel size 5, stride 1, SE enabled, activation Hard-Swish.
  • Inverted Residual Block 8 (features8): expansion factor 3, 48 to 48 channels, kernel size 5, stride 1, SE enabled, activation Hard-Swish.
  • Inverted Residual Block 9 (features9): expansion factor 6, 48 to 96 channels, kernel size 5, stride 2, SE enabled, activation Hard-Swish.
  • Inverted Residual Block 10 (features10): expansion factor 6, 96 to 96 channels, kernel size 5, stride 1, SE enabled, activation Hard-Swish.
  • Inverted Residual Block 11 (features11): expansion factor 6, 96 to 96 channels, kernel size 5, stride 1, SE enabled, activation Hard-Swish.
  • Convolutional Head (features12): 1×1 Conv2d from 96 to 576 channels, followed by BatchNorm and Hard-Swish.
  • Adaptive Average Pooling (features13): global average pool to 1×1.
  • Conv Head (features14): 1×1 Conv2d from 576 to 1024 channels, followed by Hard-Swish.
  • Flatten Layer: reshapes the 1024×1×1 tensor to a 1024-dimensional vector.
  • Fully Connected Layer (classifier0): linear 1024 to 1024, followed by Hard-Swish.
  • Dropout Layer (classifier2): dropout with $p=0.2$.
  • Fully Connected Layer (classifier3): linear 1024 to $k$ for classification.
D.2.1 Residual Neural Networks
@@ D.4 SMOTE Training
D.5 IS Training
For the classification tasks outlined in Section 4.1, we compare learn2mix and classical training with importance sampling–based neural network training (IS training) adapted from [Katharopoulos and Fleuret, 2018] and [Johansson and Lindberg, 2022]. Let $J$ denote the original training dataset, and let $\mathcal{L}_{\mathrm{ind}}^{M}(\theta^{t})\in\mathbb{R}^{M}$ denote the vector of individual cross‐entropy losses at time $t$ on a batch of size $M$ drawn uniformly from $J$. We normalize these losses to sampling probabilities, $p_{j}\in[0,1]$, sample without replacement a subset of size $b=M/2$ according to $\{p_{j}\}$, and update the model by taking a gradient step on the average loss over that subset, where:
$p_{j}=\frac{\mathcal{L}_{\mathrm{ind},j}^{M}(\theta^{t})}{\sum_{j^{\prime}=1}^{M}\mathcal{L}_{\mathrm{ind},j^{\prime}}^{M}(\theta^{t})},\quad\text{and:}\quad\mathcal{L}_{\mathrm{IS}}(\theta^{t})=\frac{1}{b}\sum_{r=1}^{b}\mathcal{L}_{\mathrm{ind},\,i_{r}}^{M}(\theta^{t}).$ (65)
In the original training dataset $J$, we use a batch size of $M$, resulting in $P=N/M$ total batches, and perform IS training on $P$ batches of size $M$ during each epoch.
D.6 CURR Training
For the classification tasks outlined in Section 4.1, we compare learn2mix and classical training with curriculum learning–based neural network training (CURR) following the self-taught scoring and fixed exponential pacing scheme of [Hacohen and Weinshall, 2019]. Let $J$ be the original training dataset, and denote by $\tilde{s}_{j}=1-\hat{p}_{j}$ the self-taught score of sample $j$, where $\hat{p}_{j}$ is the network’s confidence in the correct label after preliminary convergence training on uniform mini-batches (this warm-up stage is used only to compute $\{\tilde{s}_{j}\}$ and is not included in our reported CURR timings, nor is any analogous stage required for learn2mix). We sort the samples by increasing $\tilde{s}_{j}$ (easiest first) to obtain sorted indices $\{i_{1},\dots,i_{N}\}$. At epoch $t$, let the curriculum fraction be:
$\mathrm{frac}(t)=\min\bigl(\text{starting\_percent}\times\text{inc}^{\lfloor t/\text{step\_length}\rfloor},\,1.0\bigr),$ (66)
with $\text{starting\_percent}=0.5$, $\text{inc}=1.2$, and $\text{step\_length}=10$. We form a curriculum subset of size $\lfloor\mathrm{frac}(t)\,N\rfloor$ by taking the first $\lfloor\mathrm{frac}(t)\,N\rfloor$ sorted indices, and train on mini-batches of size $M$. The curriculum loss at time $t$ is then:
$\mathcal{L}_{\mathrm{CURR}}(\theta^{t})$ $=\frac{1}{\lfloor\mathrm{frac}(t)\,N\rfloor}\sum_{r=1}^{\lfloor\mathrm{frac}(t)\,N\rfloor}\mathcal{L}_{\mathrm{ind},\,i_{r}}^{1}(\theta^{t}),$ (67)
where $\mathcal{L}_{\mathrm{ind},\,j}^{1}(\theta^{t})\in\mathbb{R}$ is the individual cross-entropy loss on sample $j$, and each epoch processes $\lfloor\mathrm{frac}(t)\,N\rfloor/M$ batches of size $M$.
D.7 Neural Network Training Hyperparameters
The relevant hyperparameters used to train the neural networks outlined in Section D.1 are given in Table 3. All results presented in the main text were produced using these hyperparameter choices.
@@ D.7 Neural Network Training Hyperparameters
Dataset Task Optimizer
Learning
Rate ($\boldsymbol{\eta}$)
Mixing Rate ($\boldsymbol{\gamma}$)
Batch
Size ($\boldsymbol{M}$)
MNIST Classification Adam 0.0001 0.1 1000
Fashion MNIST Classification Adam 0.0001 0.5 1000
CIFAR-10 Classification Adam 7.0e-5 0.1 1000
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The paper presents a new framework for accelerating neural network convergence. We provide comprehensive empirical results and theoretical guarantees to validate this claim.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The performance gains (large or limited) afforded by learn2mix are explicitly quantified in the empirical results section, and all methods, alongside ablation studies, are thoroughly discussed in the appendix.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We provide theorems (with all relevant terms defined) in the main text, alongside comprehensive proofs in the appendix to verify the proposed theorems.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Alongside the loss function and optimizer details presented in the main text, all neural network architectures and training hyperparameters are discussed and tabulated in the appendix.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The complete code for reproducing all the empirical results for learn2mix are provided in the supplementary materials. All neural network architectures and training hyperparameters are also provided in the appendix.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The appendix details all neural network architectures utilized to generate the results presented in the main text and in the ablation studies. All hyperparameter choices are either explicitly specified in the main text/appendix, or tabulated in the appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: All empirical results in the main text and the appendix include confidence intervals in the figures and tables to explicitly declare the statistical significance of all experiments.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The compute resources (GPUs) used to produce the empirical results in the main text are specified in the empirical results section.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The presented research conforms in all aspects with the NeurIPS Code of Ethics, and the authors have reviewed the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Accelerating neural network convergence in resource constrained regimes is an important capability to ensure fast and efficient neural network training — the adoption of learn2mix can save compute cost and accelerate training. We find no negative societal impacts of our work.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: All the code is provided in the appendix and the final version will be maintained in a GitHub repository by the authors. The authors contact information will also be provided in the final version to prevent misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All relevant code and models used in the paper have been properly cited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The appendix contains comprehensive descriptions of all considered neural network architectures and modified datasets used to generate the empirical results.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.